# Emojis Decoded: Leveraging ChatGPT for Enhanced Understanding in Social Media Communications

### v2

@@ 5 Emoji Usage Intention
In addition to sentiment and meaning, previous researchers are also interested in the intention of using emojis, and ask human annotators to label emoji intentions with or without context (Hu et al. 2017; Lu et al. 2018). In this section, we follow the approaches and intention definition in the previous work (Hu et al. 2017) and ask ChatGPT to reannotate the emojis with or without any context to observe whether human and ChatGPT have similar understandings for the intention of using emojis.
@@ 5.1 Emoji Intention Annotation without Context
We first annotate emojis without any context to probe the understanding of ChatGPT on the universal intention. We select the same seven types of candidate intentions: expressing sentiment, strengthening expression, adjusting tone, expressing humor, expressing irony, expressing intimacy, and describing content, with previous work, and detailed definitions can be found in Hu et al. (2017). We select 15 emojis (5 positive, 5 neutral, and 5 negative emojis) in Hu et al. (2017) and ask ChatGPT to rate the willingness to use the emoji for each intention on a 7-point scale (7 = totally willing, 1 = not willing at all), same as the annotation method in Hu et al. (2017) and the complete prompt in Table 10.
We calculate that the average difference values between ChatGPT-rated scores and humans are -0.13 (expressing sentiment), -0.29 (strengthening expression), -0.07 (adjusting tone), -0.54 (expressing humor), -1.49 (expressing irony), -0.05 (expressing intimacy), and 1.11 (describing content), respectively. The average standard deviation (SD) of human-rated scores of seven intentions from Hu et al. (2017) is 1.643. ChatGPT and human annotators reach a high consensus on emojis in the intention of expressing sentiment, strengthening expressing, and adjusting tones. On other intentions, such as describing content, ChatGPT vary from human annotated scores but the score discrepancy is still smaller than the average human SD. Hence, we claim that the understanding of ChatGPT on the universal usage intention of emojis is generally similar to human users.
@@ 5.2 Intention of Emoji Usage in GitHub Posts
The similar intention annotation without context between ChatGPT and humans is consistent with the accurate explanation of emoji semantics and the precise annotation of emoji sentiment. When using emojis on different platforms, the intention could also vary. For example, users from GitHub may apply emojis to organize the content in the post, but seldom apply emojis to express irony. It is more challenging for ChatGPT to understand the emoji intention when given GitHub post context. In this part, we collect the same 2,000 emoji posts on GitHub with human annotations in Lu et al. (2018) and compare the annotation of emoji intention from ChatGPT and human ground truth labels. The candidate intention of emojis in GitHub posts is shown in Table 2, and the details of the definitions are shown in Lu et al. (2018).
@@
The annotated accuracy between GPT3.5 and GPT4 with human annotation is 38. 8% and 49. 0%, respectively. The detailed accuracy of GPT4 for each intention is presented in Table 2. We observe that for emojis to express sentiment and emphasize content in GitHub posts, ChatGPT gives more similar judgement with the human annotators. However, for symbol (symbolize emojis) and non-communication use (unintentional use), which is rarely applied in other social platforms, ChatGPT cannot generate the same intention annotation with annotators. To analyze the discrepancy of understanding between humans and ChatGPT, we conduct a case study and present three randomly selected posts with different annotations in Table 3.
@@
Make awesome stuff with half
the time OR make quadriply awesome
creations with the same time [Uncaptioned image]
@@
Furthermore, we analyze the ChatGPT annotations in Table 3 as well as the human labels. We feed the ChatGPT annotations back to the LLM and ask for the justification. In the first post, human annotators perceive the use of [Uncaptioned image] as an enhancement of the post’s positive sentiment, whereas ChatGPT interprets its use as a means to attract attention. In the second post, which contains the word “thanks” accompanied by [Uncaptioned image], human annotators believe that “thanks” alone conveys positive sentiment, and the emoji serves to reinforce this sentiment. In contrast, ChatGPT’s interpretation does not align with this view. The final post features [Uncaptioned image], which human annotators suggest is used for organizing content to enhance readability, while ChatGPT views it as a symbol or emphasis of a document. This case study highlights that ChatGPT’s interpretations of emoji usage in GitHub posts are not only plausible but also indicative of ChatGPT’s nuanced understanding of emoji applications.
@@
ChatGPT precisely deduces the intention of emoji usage when giving no context, and when giving a specific GitHub post, ChatGPT reaches a consensus with human annotations on the intention of emojis.
@@ 6 Emoji Usage Associated with Communities
Previous work has explored that emoji usage can be different when associated with different communities, such as platforms, languages, genders, and hashtags (Lu et al. 2018, 2016; Chen et al. 2018; Zhou and Ai 2022; Zhou et al. 2023a). In this part, we probe whether the knowledge of the associated emoji usage patterns is encoded in ChatGPT when given the prior information about the users, platform, or co-occurred hashtags. We compose the prompt by directly asking for the most used emojis given different community information, and ChatGPT outputs the most used emojis and its justification. The probing prompt for ChatGPT is designed as follows: Could you output 10 emojis that are associated with the {specific community}? where {specific community} is the associated community, such as female users, French users, hashtags, and the GitHub platform. We present the answer from ChatGPT and the findings from previous work in Tables 4, 5 and 6. In particular, emojis associated with French, female, and male users are explored in Lu et al. (2016); Chen et al. (2018) and emojis co-occurred with different hashtags and the GitHub platform are reflected in Zhou and Ai (2022); Zhou et al. (2023a).
@@
From Table 4, 5 and 6, we observe that ChatGPT depends on stereotypical associations with the attribute of the user or the platform to select the emojis and in most cases, stereotypical selection can extract the same emojis or emojis with semantics similar to the emojis shown in previous works. Both ChatGPT and existing literature identify certain emojis as indicative of gender preferences, such as [Uncaptioned image] (soccer) for male users and [Uncaptioned image] (cherry blossom) for female users. In the context of the GitHub platform, a comparison between prior findings and ChatGPT’s responses reveals a significant overlap, with seven emojis common to both lists. Furthermore, additional emojis identified by ChatGPT align with the distinctive characteristics of the GitHub environment. For instance, [Uncaptioned image] (penguin) is associated with Linux and [Uncaptioned image] (snake) is linked to Python programming.
The reliance of ChatGPT on stereotypical associations to deduce emoji usage yields insights, yet also reveals certain limitations. For instance, despite the documented preference of French users for heart-related emojis (Lu et al. 2016), the responses of ChatGPT do not reflect this trend. Furthermore, while ChatGPT employs basketball emojis in reference to #TheLastDance and Michael Jordan, it notably omits the [Uncaptioned image] (the greatest of all time) emoji, which frequently accompanies mentions of Michael Jordan. In summary, these observations suggest that although ChatGPT demonstrates an understanding of varied emoji applications across different contexts, it does not fully capture all the nuances and specificities within the domain of emoji use.
@@
ChatGPT employs stereotypical associations to infer emoji usage patterns associated with distinct communities, accurately reflecting variations in emoji use across genders, hashtags, and platforms, but fails to capture the nuances of emoji usage specific to different countries.
@@ 7 Irony Annotation with Emojis
In Section 5, ChatGPT has reached a consensus with humans about the functionality of emojis to express irony. To assess ChatGPT’s proficiency in discerning ironic emoji usage within distinct linguistic contexts, we initiated an ICL experiment on two datasets. One dataset comprises Arabic texts and annotation labels on the irony of emojis. The other consists of English tweets with emojis and annotation labels on the irony of the overall tweet. For the English dataset, we collect 405 tweets with emojis from the training dataset SemEval2018 Task 3 Subtask A (Van Hee, Lefever, and Hoste 2018), with binary annotation to indicate whether this tweet is ironic or not. Among the 405 tweets collected, 175 of them are annotated by humans as ironic, and the others are not ironic. We first use ChatGPT to predict the irony on the original tweets with emojis via ICL inference and we ask ChatGPT to annotate again by excluding the emojis in each tweet. The prompt details are in Table 12 For tweets with emojis, the accuracy of the irony prediction with GPT4 is 81.2% and when emojis are excluded, the accuracy decreases to 77.0%. Note that the accuracy drop does not mean that ChatGPT considers emojis to make irony predictions, since the irony of a sentence may change when removing emojis. We conduct a case study to show the three randomly selected irony tweets with flipped predictions after removing emojis in Table 7.
@@
To evaluate ChatGPT’s understanding of the irony in emojis for other languages, we utilize a recently published Arabic dataset with the ironic annotations: ArSarcasMoji from Hakami, Hendley, and Smith (2023). Instead of labeling the irony of sentences, ArSarcasMoji provides the ground truth irony label of emojis into three categories: sarcastic, humorous and not ironic, where the sarcastic label represents that ironic emojis play a negative sentiment role and the humorous label indicates a positive role (Hakami, Hendley, and Smith 2023). Since the label distribution in the original dataset is biased, we first downsample sentences with the label “not ironic” and “sarcastic” to make a balanced dataset for evaluation. We compose the prompt with instruction, the detailed definition of “humorous” and “saracastic”, and the examples of Arabic sentences with labels (the complete prompt in Table 13). The prediction accuracy for GPT4 and GPT3.5 is 60.1% and 63.1%, respectively, indicating that ChatGPT can understand the functionality of emojis to express irony in the Arabic context.
@@
In the emoji irony classification task, ChatGPT has promising performance in distinguishing whether the emoji is ironic or not and in the tweet irony classification task, ChatGPT considers the existence of emoji to decide the overall irony of a tweet.
@@ 8 Emoji Recommendation
After knowing that ChatGPT can precisely point out the sentiments, meanings, and intentions of emojis, our next question is whether ChatGPT relies on its knowledge to recommend suitable emojis.
@@ 8.1 Recommendation Based on Text Context
For the experiment, we choose the SemEval 2018 Task 2 with English and Spanish sets for emoji recommendation (Barbieri et al. 2018). We select the validation sets with 5,000 English and Spanish tweets for inference. The task pre-defines a list of 20 or 19 emojis for models to choose for English and Spanish tasks, respectively, and our prompt is composed of the English instruction including the predefined emoji candidates and five exemplars from the English or Spanish training set, depending on the prediction text. The prompt details are shown in Table 14.
For emoji recommendation on English tweets, GPT4 and GPT3.5 can achieve 22.12% and 21.52% accuracy and the finetuned FastText classifier can achieve 42.56% accuracy. On Spanish tweets, GPT4 and GPT3.5 can obtain 20.14% and 20.96% accuracy while the finetuned FastText can obtain 31.63% accuracy. These results suggest that, on a closed-set emoji recommendation, the ICL of ChatGPT cannot beat the traditional fine-tuned classifer, but ChatGPT’s performance is consistent across the languages while the fine-tuned classifier meets a large performance drop. We can observe the hallucinations of ChatGPT when performing the emoji recommendation, and 6.1% and 3.7% of tweets are annotated with the emoji not in the pre-defined candidates for GPT3.5 and GPT4 on English tweets.
@@ 8.2 Recommendation with User Identity
Previous work suggests the different emoji preference depending on the user identity (Chen et al. 2018; Lu et al. 2016) so when performing the emoji recommendation, it is also important to consider the user demographic attributes to recommend the proper emoji choices. ChatGPT has been shown to have the potential to play as a communicative agent with different roles (Li et al. 2023a). Moreover, as shown in Section 6, the knowledge about the discrepancy in the emoji usage patterns of different communities is encoded in ChatGPT. In this part, we investigate whether ChatGPT has the ability to consider user identity when recommending the emoji to the tweet.
We ask ChatGPT to play as a female or male Twitter user to give the recommendation of the emojis, respectively, and add another sentence at the beginning of the prompt like the previous work (Li et al. 2023a): I want you to act as a female/male Twitter user. Never forget your role. We repeat the GPT4 ICL experiment on 5,000 English tweets. Acting as a male or female twitter user can obtain 21.04% and 20.94% accuracy, respectively, and 1,110 tweets have different predictions. We present the distribution of the top 10 predicted emojis for female or male users in the first two rows of Table 8.
@@
An examination of the first two rows of Table 8 reveals a comparable emoji distribution with two notable exceptions: a higher frequency of [Uncaptioned image] (two hearts) when representing female users, and a higher usage of [Uncaptioned image] in representations of male users. This finding is consistent with ChatGPT’s understanding of usage differences for female and male users in Section 6 and the findings from previous work (Chen et al. 2018). To further explore ChatGPT’s comprehension of variations in emoji usage across different countries, an aspect not revealed by the results in Section 6, we conducted additional experiments. In these, we simulate interactions while portraying users from France and the United States, respectively. When asking ChatGPT to act as the French user, we translate the prompts and tweets to annotate in the French language. We also present the predicted emoji distribution in the last two rows of Table 8.
From the last two rows, we can observe a similar emoji distribution for ChatGPT to portray users from France or the USA. In conjunction with the findings presented in Section 6 regarding ChatGPT’s interpretation of emoji usage by French users, it can be concluded that ChatGPT lacks awareness of the variations in emoji usage across different countries.
@@
ChatGPT does not outperform the fine-tuned traditional classifier on the closed-set emoji recommendation task, but ChatGPT can depend on user’s identity to make more personalized emoji recommendations.

### v3

@@ 5 Emoji Usage Intention
In addition to sentiment and meaning, previous researchers have also explored the intention behind using emojis, asking human annotators to label emoji intentions with or without context (Hu et al. 2017; Lu et al. 2018). In this section, we follow the approaches and intention definitions from previous work (Hu et al. 2017) and ask ChatGPT to reannotate the emojis with or without context. This allows us to observe whether humans and ChatGPT have similar understandings of the intention behind using emojis.
@@ 5.1 Emoji Intention Annotation without Context
First, we annotate emojis without any context to assess ChatGPT’s understanding of their universal intentions. We use the same seven types of candidate intentions as in previous work: expressing sentiment, strengthening expression, adjusting tone, expressing humor, expressing irony, expressing intimacy, and describing content. Detailed definitions of these intentions can be found in Hu et al. (2017). We select 15 emojis (5 positive, 5 neutral, and 5 negative) from Hu et al. (2017) and ask ChatGPT to rate the willingness to use each emoji for each intention on a 7-point scale (7 = totally willing, 1 = not willing at all), following the same annotation method as in Hu et al. (2017). The complete prompt is shown in Table 13 in Appendix.
We calculate the average difference between ChatGPT-rated scores and human ratings as follows: -0.13 (expressing sentiment), -0.29 (strengthening expression), -0.07 (adjusting tone), -0.54 (expressing humor), -1.49 (expressing irony), -0.05 (expressing intimacy), and 1.11 (describing content). Given that the average standard deviation (SD) of human-rated scores is 1.643 (Hu et al. 2017), ChatGPT and human annotators reach a high consensus on the intentions of expressing sentiment, strengthening expression, and adjusting tone. For other intentions, such as describing content, ChatGPT’s scores differ from human ratings, but the discrepancy is smaller than the average human SD. Hence, we conclude that ChatGPT’s understanding of the universal usage intention of emojis is generally similar to that of human users.
@@ 5.2 Intention of Emoji Usage in GitHub Posts
The similar intention annotation without context between ChatGPT and humans aligns with the accurate explanation of emoji semantics and the precise annotation of emoji sentiment. However, the intention behind using emojis can vary across different platforms. For example, users on GitHub may use emojis to organize content within a post but rarely to express irony. Understanding emoji intention in GitHub posts poses a greater challenge for ChatGPT.
In this part, we collect the same 2,000 emoji posts on GitHub with human annotations from Lu et al. (2018) and compare ChatGPT’s annotations of emoji intentions with the human ground truth labels. The candidate intentions of emojis in GitHub posts are shown in Table 2, with detailed definitions provided in Lu et al. (2018).
@@
The annotated accuracy between GPT3.5 and GPT4 compared with human annotations is 38.8% and 49.0%, respectively. The detailed accuracy of GPT4 for each intention is presented in Table 2. We observe that for emojis used to express sentiment and emphasize content in GitHub posts, ChatGPT’s judgments are more aligned with human annotators. However, for symbols (symbolize emojis) and non-communication use (unintentional use), which are rarely applied on other social platforms, ChatGPT’s intention annotations do not match those of human annotators.
We also conduct a detailed case study and present three randomly selected posts with different annotations in Table 3 to further analyze the extent to which ChatGPT’s responses align with human understanding.
@@
Make awesome stuff
with half the time OR
make quadriply
awesome creations
with the same time [Uncaptioned image]
@@
Furthermore, we feed the ChatGPT annotations in Table 3 back to the LLM and ask for justification. In the first post, human annotators perceive the use of [Uncaptioned image] as an enhancement of the post’s positive sentiment, whereas ChatGPT interprets its use as a means to attract attention. In the second post, which contains the word “thanks” accompanied by [Uncaptioned image], human annotators believe that “thanks” alone conveys positive sentiment, and the emoji serves to reinforce this sentiment. ChatGPT’s interpretation also notices that the emoji conveys the sentiment and gives a similar annotation: sentiment expressed. The final post features [Uncaptioned image], which human annotators suggest is used for organizing content to enhance readability, while ChatGPT views it as a symbol or emphasis of a document. From the case study, we find that even in incorrect annotations, ChatGPT still captures the primary intention behind the emoji usage. This case study highlights that ChatGPT’s interpretations of emoji usage in GitHub posts are not only plausible but also indicative of ChatGPT’s nuanced understanding of emoji applications.
@@
ChatGPT precisely deduces the intention of emoji usage without context and, when provided with a specific GitHub post, reaches a consensus with human annotations on the intention of emojis.
@@ 6 Emoji Usage Associated with Communities
While ChatGPT has demonstrated a deep understanding of emoji semantics, intentions, and sentiments, previous work suggests that emoji usage can vary significantly across different communities, such as platforms, languages, genders, and hashtags (Lu et al. 2018, 2016; Chen et al. 2018; Zhou and Ai 2022; Zhou et al. 2023a). In this section, we investigate whether ChatGPT encodes knowledge of these associated emoji usage patterns when provided with prior information about users, platforms, or co-occurring hashtags. We construct the prompt by directly asking for the most used emojis given different community information, and ChatGPT outputs the most used emojis along with its justification. The probing prompt for ChatGPT is designed as follows:
Could you output 10 emojis that are associated with the specific community?
where specific community refers to the associated community, such as female users, French users, specific hashtags, or the GitHub platform.
We present ChatGPT’s answers and compare them with findings from previous work in Tables 4, 5, and 6. In particular, emojis associated with French, female, and male users are explored in Lu et al. (2016); Chen et al. (2018), and emojis co-occurring with different hashtags and on the GitHub platform are reflected in Zhou and Ai (2022); Zhou et al. (2023a); Lu et al. (2018).
@@
From Tables 4, 5, and 6, we observe that ChatGPT relies on stereotypical associations with user attributes or platform characteristics to select emojis. In most cases, this stereotypical selection results in choosing the same or semantically similar emojis to those identified in previous works.
For example, both ChatGPT and previous works mention [Uncaptioned image] (soccer) and [Uncaptioned image] (cherry blossom) as emojis associated with male and female users, respectively. Regarding the most frequently used emojis on the GitHub platform, 7 emojis overlap between previous findings and ChatGPT’s answers. Additionally, the other emojis from ChatGPT also reflect attributes of the GitHub platform, such as [Uncaptioned image] (penguin) for Linux and [Uncaptioned image] (snake) for Python.
However, since ChatGPT utilizes stereotypical associations to infer associated emojis, the results may sometimes be incomplete. For example, while previous research indicates that French users often use heart-related emojis (Lu et al. 2016), this is not reflected in ChatGPT’s responses. Similarly, although ChatGPT uses basketball emojis to indicate #TheLastDance about Michael Jordan, it does not include [Uncaptioned image] (the greatest of all time), which often co-occurs with references to Michael Jordan. In summary, the results indicate that ChatGPT notices differences in emoji usage across different contexts, but it cannot cover all facts or details in the emoji domain.
@@
ChatGPT employs stereotypical associations to infer emoji usage patterns associated with distinct communities, accurately reflecting variations in emoji use across genders, hashtags, and platforms. However, it fails to capture the nuances of emoji usage specific to different countries.
@@ 7 Irony Annotation with Emojis
In Section 5, ChatGPT has reached a consensus with humans about the functionality of emojis to express irony. To assess ChatGPT’s proficiency in discerning ironic emoji usage within distinct linguistic contexts, we initiated an ICL experiment on two datasets. One dataset comprises Arabic texts with annotation labels on the irony of emojis. The other consists of English tweets with emojis and annotation labels on the irony of the overall tweet.
For the English dataset, we collected 405 tweets with emojis from the training dataset SemEval2018 Task 3 Subtask A (Van Hee, Lefever, and Hoste 2018), which includes binary annotations indicating whether each tweet is ironic or not. Among the 405 tweets collected, 175 are annotated by humans as ironic, and the remaining are not.
We first used ChatGPT to predict the irony in the original tweets with emojis via ICL inference. We then asked ChatGPT to annotate the tweets again, this time excluding the emojis. The prompt details are provided in Table 15. For tweets with emojis, the accuracy of irony prediction with GPT4 is 81.2%. When the emojis are excluded, the accuracy decreases to 77.0%. Note that the accuracy drop does not necessarily mean that ChatGPT uses emojis to make irony predictions, as the irony of a sentence may change when emojis are removed. We conducted a case study and present three randomly selected ironic tweets with flipped predictions after removing emojis in Table 7.
@@
To evaluate ChatGPT’s understanding of irony in emojis for other languages, we utilize a recently published Arabic dataset with ironic annotations: ArSarcasMoji from Hakami, Hendley, and Smith (2023). Instead of labeling the irony of sentences, ArSarcasMoji provides the ground truth irony label of emojis in three categories: sarcastic, humorous, and not ironic. The sarcastic label indicates that ironic emojis play a negative sentiment role, while the humorous label indicates a positive role (Hakami, Hendley, and Smith 2023).
Since the label distribution in the original dataset is biased, we first downsample sentences labeled as “not ironic” and “sarcastic” to create a balanced dataset for evaluation. We compose the prompt with instructions, detailed definitions of “humorous” and “sarcastic,” and examples of Arabic sentences with labels (the complete prompt is provided in Table 16). The prediction accuracy for GPT4 and GPT3.5 is 60.1% and 63.1%, respectively, indicating that ChatGPT can understand the functionality of emojis to express irony in the Arabic context.
@@
In the emoji irony classification task, ChatGPT demonstrates promising performance in distinguishing whether an emoji is ironic or not. In the tweet irony classification task, ChatGPT considers the presence of emojis when determining the overall irony of a tweet.
@@ 8 Emoji Recommendation
Emoji recommendation, a well-defined and widely adopted task, is essential for evaluating models’ understanding of emoji usage (Barbieri et al. 2020; Lyu et al. 2024). This task requires models to grasp both the semantics of emojis and their application context. After establishing that ChatGPT can identify the sentiments, meanings, and intentions behind emojis, we examine its ability to recommend suitable emojis based on a given text context.
@@ 8.1 Recommendation Based on Text Context
For the experiment, we use the SemEval 2018 Task 2 dataset for emoji recommendation in English and Spanish (Barbieri et al. 2018). We select the validation sets with 5,000 English and Spanish tweets for inference. The task pre-defines a list of 20 emojis for English and 19 for Spanish, and our prompt includes the predefined emoji candidates and five exemplars from the respective training set. The prompt details are shown in Table 17.
For emoji recommendation on English tweets, GPT4 and GPT3.5 achieve 22.12% and 21.52% accuracy, respectively, while the fine-tuned FastText classifier achieves 42.56% accuracy. For Spanish tweets, GPT4 and GPT3.5 obtain 20.14% and 20.96% accuracy, respectively, while the fine-tuned FastText classifier achieves 31.63% accuracy. These results suggest that, in a closed-set emoji recommendation, the ICL of ChatGPT cannot outperform the traditional fine-tuned classifier. However, ChatGPT’s performance is consistent across languages, whereas the fine-tuned classifier experiences a significant performance drop. Furthermore, we observe ChatGPT hallucinations during emoji recommendation, with 6.1% and 3.7% of English tweets being annotated with emojis not in the predefined candidates for GPT3.5 and GPT4, respectively.
@@ 8.2 Recommendation with User Identity
Previous work suggests that emoji preferences differ based on user identity (Chen et al. 2018; Lu et al. 2016), so it is important to consider user demographic attributes when recommending emojis. ChatGPT has shown the potential to act as a communicative agent with different roles (Li et al. 2023a). Moreover, as shown in Section 6, ChatGPT encodes knowledge about discrepancies in emoji usage patterns among different communities. In this part, we investigate whether ChatGPT can consider user identity when recommending emojis for tweets.
We ask ChatGPT to act as a female or male Twitter user and provide emoji recommendations accordingly. We add another sentence at the beginning of the prompt, as in previous work (Li et al. 2023a): I want you to act as a female/male Twitter user. Never forget your role. We repeat the GPT4 ICL experiment on the same 5,000 English tweets. Acting as a male or female Twitter user, ChatGPT achieves 21.04% and 20.94% accuracy, respectively, with 1,110 tweets having different predictions. Although the provided gender identity does not significantly change the accuracy, it altered more than 20% of the predictions of the recommended emojis. We also performed a Chi-Square Test (Rana and Singhal 2015) to examine the significance of the difference in emoji distributions between the two prompts, yielding a p-value of less than 0.01. We present the distribution of the top 10 predicted emojis for female and male users in the first two rows of Table 8.
@@
An examination of the first two rows of Table 8 reveals a comparable emoji distribution with two notable exceptions: a higher frequency of [Uncaptioned image] (two hearts) when representing female users, and a higher usage of [Uncaptioned image] (smiling face with sunglasses) in representations of male users. This finding is consistent with ChatGPT’s understanding of usage differences for female and male users in Section 6 and the findings from previous work (Chen et al. 2018).
To further explore ChatGPT’s comprehension of variations in emoji usage across different countries, an aspect not revealed by the results in Section 6, we conducted additional experiments. In these, we simulate interactions while portraying users from France and the United States, respectively. When asking ChatGPT to act as a French user, we translate the prompts and tweets to French. We present the predicted emoji distribution in the last two rows of Table 8.
From the last two rows, although the p-value of Chi-Square Test is still smaller than 0.01, we cannot find a significant change in the occurrence numbers for a specific emoji when portraying users from France or the USA. In conjunction with the findings presented in Section 6 regarding ChatGPT’s interpretation of emoji usage by French users, it can be concluded that ChatGPT lacks awareness of the variations in emoji usage across different countries.
@@
ChatGPT does not outperform the fine-tuned traditional classifier on the closed-set emoji recommendation task. However, when provided with user gender information, ChatGPT can alter its emoji recommendations.